# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

### v2

@@ 4.1.2 Values
Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group [Schwartz, 1992]. Values offer a powerful lens for understanding LLM behavior. For example, Ye et al. [2025a] show how different values contribute to the safety of LLMs; Liu et al. [2024b] reveal that different spiritual values affect LLMs in social-fairness scenarios; and Sorensen et al. [2024b] demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
A growing body of research has applied diverse value theories and instruments to assess the value orientations of LLMs. Schwartz’s Value Theory, with its ten basic values and higher-order dimensions, is the most widely used, often via the Schwartz Value Survey (SVS) or Portrait Values Questionnaire (PVQ) [Schwartz, 1992, Schwartz et al., 2001]. Adaptations and contextualized prompts are increasingly common to better suit LLMs’ operational context [Ren et al., 2024, Shen et al., 2024]. The World Values Survey (WVS) has also been used to probe LLMs’ alignment with societal-level value dimensions [Haerpfer et al., 2022, Kim and Baek, 2024]. Cross-cultural frameworks like Hofstede’s Values Survey Module (VSM) and the GLOBE study extend this analysis to workplace and leadership-related cultural dimensions, with both direct and adapted inventories employed [Hofstede, 1984, House et al., 2004, Ren et al., 2024]. Social Value Orientation (SVO) frameworks, using tools like the SVO Slider Measure, focus on LLMs’ prosocial versus proself tendencies [Murphy et al., 2011, Zhang et al., 2024c]. Additionally, researchers have developed localized or custom inventories to capture region-specific or topical values, and some propose novel, bottom-up value taxonomies for LLMs using psycholexical data [Meadows et al., 2024, Xu et al., 2023, Biedma et al., 2024, Ye et al., 2025b]. Collectively, these studies reveal both the adaptability of human value frameworks to LLMs and the need for tailored approaches.
@@
Research indicates that LLMs display distinct and systematic value patterns. According to Schwartz’s Value Theory, LLMs tend to prioritize Self-Transcendence and Conservation. They exhibit stronger inclinations toward Universalism, Benevolence, Conformity, and Security, while opposing Power and Achievement [Rozen et al., 2024, Zhang et al., 2023a, Hadar-Shoval et al., 2024, Fischer et al., 2023]. WVS surveys further suggest that LLMs generally prefer Self-Expression values over Survival values [Chiu et al., 2025]. Studies using the VSM and GLOBE frameworks emphasize LLMs’ strong focus on Humane and Performance Orientation, with moderate Assertiveness [Li et al., 2024b]. Additionally, when assessed through SVO, advanced LLMs predominantly show Prosocial tendencies [Zhang et al., 2024c].
Different models exhibit varied value orientations [Kovač et al., 2024, Chiu et al., 2025, Xu et al., 2023, Li et al., 2024b, Duan et al., 2023]. Versions within the same model family show evolutionary trends in values, potentially influenced by safety alignment, capability advancement, and shifting societal expectations [Duan et al., 2023, Moore et al., 2024, Kim and Baek, 2024]. Generally, larger models align more closely with desirable human values [Shen et al., 2024, Jiang et al., 2024a, Kim and Baek, 2024, Zhong et al., 2024].
Cross-cultural research suggests that LLMs may embody a blend of cultural values, integrating perspectives from diverse backgrounds [Kovač et al., 2023]. However, they generally exhibit a tendency toward Western liberal values [Kim and Baek, 2024]. LLMs can display different values based on profiling prompts [Kharchenko et al., 2024, Zhong et al., 2024, Karinshak et al., 2024]. The context-dependency of their values is concerning and challenges the stability assumptions in human value theories [Shen et al., 2024, Meadows et al., 2024, Moore et al., 2024, Chiu et al., 2025, Xu et al., 2023, Kovač et al., 2023].
Comparing value theories and suggesting use cases.
Schwartz emphasizes universal human motivations; WVS examines societal-level value shifts; VSM and GLOBE address cross-cultural dimensions, with GLOBE emphasizing leadership. SVO, in contrast, focuses on individual-level social preferences, i.e., prosocial vs. proself. While Schwartz and SVO overlap in analyzing individual motivations, the former is broader, and the latter narrowly focuses on decision-making in resource allocation. WVS, VSM, and GLOBE share cultural-level dimensions but differ in focus—WVS on modernization, VSM on workplace-related cultural differences, and GLOBE on leadership and organizational practices. For evaluating individual values in LLMs, use Schwartz for a comprehensive understanding of universal motivations or SVO for interpersonal decision-making studies. In contrast, choose WVS for evaluating broad societal values in LLMs. When evaluating cultural dimensions in LLMs, use VSM for workplace cultural differences or GLOBE for leadership and organizational practices. In addition, Schwartz’s theory bridges individual and societal-level values, making it versatile for mixed-focus studies.
@@ 4.1.3 Morality
Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong [Long and Sedley, 1987]. It is crucial to conduct moral assessments of LLMs to ensure their ethical deployment. A large body of research applies the Moral Foundations Theory (MFT) [Graham et al., 2009], primarily through the Moral Foundations Vignettes (MFVs) [Clifford et al., 2015], the Moral Foundations Questionnaire (MFQ) [Graham et al., 2009], the MFQ-2 [Atari et al., 2023], and the Moral Foundations Dictionary (MFD) [Graham et al., 2009]. These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [Tlaie, 2024, Nunes et al., 2024, Abdulhai et al., 2024], persona-driven exploration [Münker, 2024], or evaluating internal moral coherence [Nunes et al., 2024]. Other prominent moral theories and instruments include Kohlberg’s Theory via the Defining Issues Test (DIT) [Kohlberg, 1964], the Consequentialist-Deontological distinction [Beauchamp, 2001], the PEW 2013 Global Attitudes Survey [Center, 2013], and localized moral theories [Liu et al., 2024d, Takeshita et al., 2023]. Specialized datasets are also curated for scalable and comprehensive morality evaluation [Hendrycks et al., 2021, Jinnai, 2024, Marraffini et al., 2024, Jin et al., 2024b]
@@
LLMs are generally characterized by a rationalist and consequentialist focus, often prioritizing harm minimization and fairness [Neuman et al., 2025]. Despite that, they show divergence in ethical reasoning [Neuman et al., 2025] and moral preferences [Tanmay et al., 2023, Jin et al., 2024b, Bonagiri et al., 2024, Meijer et al., 2024]. In some aspects, most LLMs align with human moral standards [Tanmay et al., 2023, Nunes et al., 2024, Takemoto, 2024], which may be attributed to their extensive exposure to conventional ethical values during training. On the other hand, some research presents a less optimistic view, uncovering significant deviations of LLMs from human moral preferences [Marraffini et al., 2024, Vida et al., 2024, Ahmad and Takemoto, 2024]. Regarding the underlying mechanisms of LLMs’ moral reasoning, Nunes et al. [2024], Ji et al. [2024], Simmons [2023] suggest that LLMs primarily exhibit imitation rather than genuine conceptual understanding.
Comparing moral theories/instruments and suggesting use cases.
Instruments based on MFT are well-suited for measuring LLMs’ alignment with intuitive moral principles across cultural or political dimensions. Tools like MFQ, MFQ-2, MFVs, and MFD are particularly effective when investigating ideological leanings, bias, or internal consistency in abstract vs. contextual moral judgments [Nunes et al., 2024]. In contrast, Kohlberg’s DIT explores structured moral reasoning, offering a developmental perspective on how LLMs process ethical dilemmas [Khandelwal et al., 2024, Tanmay et al., 2023]. DIT is valuable for longitudinal studies and evaluating whether models show higher-level moral reasoning (e.g., post-conventional reasoning). Frameworks such as the Consequentialist-Deontological distinction and global surveys like the PEW 2013 study are useful for categorizing ethical tendencies or analyzing cultural homogeneity [Neuman et al., 2025, Meijer et al., 2024]. Localized tools (e.g., CMD, JCM) help assess cultural alignment [Liu et al., 2024d, Takeshita et al., 2023]. Meanwhile, datasets like ETHICS provide morally charged scenarios for detecting normative behavior and fine-tuning LLM responses in high-stakes ethical domains [Hendrycks et al., 2021, Jin et al., 2024b, Marraffini et al., 2024].
@@ 4.1.4 Attitudes and opinions
Attitudes are always attitudes about something. This implies three necessary elements: first, there is the object of thought, which is both constructed and evaluated. Second, there are acts of construction and evaluation. Third, there is the agent, who is doing the constructing and evaluating. We can therefore suggest that, at its most general, an attitude is the cognitive construction and affective evaluation of an attitude object by an agent [Bergman, 1998]. We use the term "attitude" to encompass both attitudes and opinions, following [Bergman, 1998, Ma et al., 2024a]. Most research on LLM attitudes examines political attitudes and public opinions [Ma et al., 2024a], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [Durmus et al., 2023, Santurkar et al., 2023, Sanders et al., 2023, Lee et al., 2024c, Hartmann et al., 2023]. For a broader discussion on bias and fairness in LLMs, we refer readers to [Gallegos et al., 2024, Ranjan et al., 2024].
Researchers assess LLMs’ political attitudes using standardized questionnaires and scales from political science and social psychology. US-focused instruments include the American National Election Studies (ANES) [Qi et al., 2024] and the American Trends Panel (ATP) [Santurkar et al., 2023]. For cross-national comparisons, studies use tools like the German Longitudinal Election Study (GLES) [Ma et al., 2024b, Ball et al., 2025, von der Heyde et al., 2024]. The Political Compass Test (PCT) is also widely used to map LLMs within multidimensional political spectrums [Azzopardi and Moshfeghi, 2024, Bernardelle et al., 2024, Hartmann et al., 2023, Röttger et al., 2024].
Other survey instruments include the General Social Survey (GSS) [Kim and Lee, 2023], the American Community Survey (ACS) [Dominguez-Olmedo et al., 2024], the Canadian Election Study (CES) [Sanders et al., 2023], the European Social Survey (ESS) [Geng et al., 2024], the Survey of Russian Elites [Kalinin, 2023], and the Supreme Court Case Political Evaluation (SCOPE) [Xu et al., 2025c]. Additionally, researchers have developed specialized datasets and tailored tools for LLMs, such as OpinionQA [Santurkar et al., 2023], a comprehensive dataset based on ATP surveys; IssueBench [Röttger et al., 2025], a benchmark covering controversial issues with multiple response formats; GlobalOpinionQA [Durmus et al., 2023], which extends political opinion evaluation to cross-cultural contexts; and OpinionGPT [Haller et al., 2023], a web tool that demonstrates how input data biases influence model outputs.
@@
Pretraining data inherently contains socially biased opinions and perspectives, which can amplify political polarization in LLMs [Feng et al., 2023, Xu et al., 2025c]. Most studies identify a misalignment between LLM outputs and human opinions [Santurkar et al., 2023, von der Heyde et al., 2024, Yang et al., 2024, Dormuth et al., 2025], with many concluding that LLMs exhibit a left-leaning political bias [Rozado, 2023, Hartmann et al., 2023, Ceron et al., 2024, Bernardelle et al., 2024, Ma et al., 2024b]. Cross-cultural comparative studies further reveal Western-centric tendencies, demonstrating limited understanding of non-English political perspectives or multi-partisan systems [Qi et al., 2024]. These findings suggest structural limitations in the models’ political understanding. It is also shown that the degree and manifestation of bias vary significantly across contexts and domains. For example, political-electoral propositions exhibit different bias patterns than socioeconomic issues like climate change [Wu et al., 2023]. In contrast, some researchers offer more optimistic perspectives, emphasizing LLMs’ potential to simulate population-level opinions and supplement traditional survey methods [Dominguez-Olmedo et al., 2024, Sanders et al., 2023, Kalinin, 2023, Wu et al., 2023, Jiang et al., 2024d, Argyle et al., 2023, Bisbee et al., 2024, Sun et al., 2024]. They suggest that with appropriate prompt design, calibration methods, and fine-tuning, LLMs can generate opinion distributions closely approximating human group distributions [Jiang et al., 2022, Wu et al., 2023]. Such advancements indicate promising directions for using LLMs in opinion research, though their current limitations and biases must still be acknowledged.
Comparing attitude measurement tools and suggesting use cases.
The ANES offers the most comprehensive framework for American political contexts, making it ideal for studying model performance on U.S. social issues [Jiang et al., 2022]. Cross-national surveys and datasets such as the ESS, GLES, and GlobalOpinionQA provide broader cultural perspectives across diverse sociopolitical environments [Geng et al., 2024, Ball et al., 2025, Ma et al., 2024b, von der Heyde et al., 2024, Durmus et al., 2023]. The ATP is better suited for researching model responses to social controversies [Santurkar et al., 2023], while the PCT serves as a theory-driven political spectrum measurement tool, offering simplified yet intuitive insights into LLM political positioning [Röttger et al., 2024]. Other surveys have distinct regional and sociopolitical focuses: the CES examines political participation in North American multicultural democracies [Sanders et al., 2023]; Survey of Russian Elites provides insights into non-Western political perspectives [Kalinin, 2023]; the SCOPE delivers in-depth assessments of judicial attitudes [Xu et al., 2025c]; and the ACS serves as a tool for researching socioeconomic issues and group representation [Dominguez-Olmedo et al., 2024].
@@ 4.2 Measuring cognitive constructs
Traditional NLP benchmarks are insufficient to capture and analyze emergent cognitive constructs [Hagendorff et al., 2024, Ying et al., 2025a]. To address this gap, researchers are adapting related psychometric techniques to evaluate LLM abilities. Hagendorff et al. [2024] introduce the concept of machine psychology, reviewing emergent LLM capabilities and advocating for psychometric evaluation. They categorize these cognitive constructs into four key aspects: heuristics and biases, social interactions, psychology of language, and learning and cognitive capabilities. Structured based on this taxonomy, Table 4 summarizes the representative cognitive constructs measured in recent research.
Aspects Sub-aspects Related work
Heuristics & biases - Yax et al. [2024], Sundaram and Alwar [2024], Momennejad et al. [2023], Echterhoff et al. [2024], Saeedi et al. [2025], Ranaldi and Zanzotto [2024], Macmillan-Scott and Musolesi [2024], Ando et al. [2023], Malberg et al. [2024], Xie et al. [2024b], Hayes et al. [2024], Shah et al. [2024], Chen and Eger [2025], Tang and Kejriwal [2024], Talboy and Fuller [2023], Coda-Forno et al. [2024], Abramski et al. [2023], Binz and Schulz [2023], Xie et al. [2024a], Kumar et al. [2024], Schulze Buschoff et al. [2025], Scholten et al. [2024], Shaikh et al. [2024], Thorstad [2023], Hagendorff et al. [2023], Healey et al. [2024], Bai et al. [2024], Castello et al. [2024]
Social interactions Theory of Mind Sarıtaş et al. [2025], Ma et al. [2023], Mao et al. [2024b] (reviews); Ullman [2023], Sap et al. [2022], Kosinski [2023a], Shapira et al. [2023a], Gandhi et al. [2023], Strachan et al. [2024a], Kosinski [2023b], Moghaddam and Honey [2023], Street et al. [2024], Sclar et al. [2023], van Duijn et al. [2023], He et al. [2023], Holterman and Deemter [2023], Chen et al. [2024d], Zhu et al. [2024c], Hou et al. [2024], Amirizaniani et al. [2024], Soubki et al. [2024], Jones et al. [2024], Lin et al. [2024], Sclar et al. [2024], Sadhu et al. [2024], Lorè et al. [2024], Sarangi et al. [2025], Yu et al. [2025], Tang and Belle [2024], Kim et al. [2025a], Chan et al. [2024], Nickel et al. [2024], Xu et al. [2024a], Leer et al. [2023], Jamali et al. [2023], Li et al. [2023b], Wilf et al. [2023], Chen et al. [2024c], Strachan et al. [2024b], Xu et al. [2025a], Zhang et al. [2025], Shinoda et al. [2025], Jin et al. [2024a], Yang et al. [2025a], Zhou et al. [2023a], Pi et al. [2024], Riemer et al. [2025], Tan et al. [2024].
Emotional Intelligence Sorin et al. [2024], Raj [2024] (reviews); Sabour et al. [2024], Paech [2023], Wang et al. [2023b], Elyoseph et al. [2023], Patel and Fan [2023], Vzorinab et al. [2024], Hu et al. [2025], Li et al. [2024e], Zhao et al. [2024], Chen et al. [2024b], Huang et al. [2024b], Schaaff et al. [2023], Li et al. [2023a], Huang et al. [2023b], Lee et al. [2024f], Welivita and Pu [2024]
Social Intelligence Leng and Yuan [2023], Mittelstädt et al. [2024], Shapira et al. [2023b], Zhou et al. [2024b], Wang et al. [2024d], Liu et al. [2024c], Wang et al. [2024a], Mou et al. [2024], Liang et al. [2024], Zadeh et al. [2019], Xu et al. [2024b], Guo et al. [2023a], Kovač et al. [2024, 2021], Mathur et al. [2024], Schoenegger et al. [2025]
Psychology of language Language comprehension Chang and Bergen [2024] (reviews); Duan et al. [2024a], Wang et al. [2024b], Wilcox et al. [2021], Arehalli et al. [2022], Li et al. [2024a], Steuer et al. [2023], Duan et al. [2024b], Qiu et al. [2024], Bojić et al. [2023], Hu et al. [2023], Ruis et al. [2023], Lee et al. [2024b], Seals and Shalin [2023], Hu and Levy [2023], He et al. [2024], Sun and Wang [2024], Hong et al. [2023], Amouyal et al. [2024], Ide et al. [2024], Lee et al. [2024b], Huff and Ulakçi [2024], Zhou et al. [2023b], Tian et al. [2024]
Language generation Seals and Shalin [2023], Tian et al. [2024], Tang and Kejriwal [2024], Stevenson et al. [2022], Hubert et al. [2023], Chakrabarty et al. [2024], Orwig et al. [2024], Lee and Chung [2024], Boussioux et al. [2024], Miaschi et al. [2024], Bellemare-Pepin et al. [2024], Cai et al. [2024]
Language acquisition Steuer et al. [2023], Shah et al. [2024], Frank [2023b], Riva et al. [2024]
Learning and cognitive capabilities - Niu et al. [2024] (reviews); Galatzer-Levy et al. [2024], Wang et al. [2024e], Dayan et al. [2024], Wu et al. [2025], Sartori and Orrú [2023], Webb et al. [2022], Song et al. [2024a], Zhang et al. [2023b], Wang et al. [2024e], Coda-Forno et al. [2024], Zeng et al. [2024c], Lv et al. [2024]
Table 4: Representative LLM cognitive constructs in LLM psychometrics.

### v3

@@ 4.1.2 Values
Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group [schwartz1992universals]. Values offer a powerful lens for understanding LLM behavior. For example, ye2025gpv show how different values contribute to the safety of LLMs; liu2024measuring reveal that different spiritual values affect LLMs in social-fairness scenarios; and sorensen2024position demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
A growing body of research has applied diverse value theories and instruments to assess the value orientations of LLMs. Schwartz’s Value Theory, with its ten basic values and higher-order dimensions, is the most widely used, often via the Schwartz Value Survey (SVS) or Portrait Values Questionnaire (PVQ) [schwartz1992universals, schwartz2001extending]. Adaptations and contextualized prompts are increasingly common to better suit LLMs’ operational context [ren2024valuebench, shen2024valuecompass]. The World Values Survey (WVS) has also been used to probe LLMs’ alignment with societal-level value dimensions [haerpfer2022wvs, kim2024exploring]. Cross-cultural frameworks like Hofstede’s Values Survey Module (VSM) and the GLOBE study extend this analysis to workplace and leadership-related cultural dimensions, with both direct and adapted inventories employed [hofstede1984culture, house2004globe]. Social Value Orientation (SVO) frameworks, using tools like the SVO Slider Measure, focus on LLMs’ prosocial versus proself tendencies [murphy2011svo_sliders, zhang2024heterogeneous]. Additionally, researchers have developed localized or custom inventories to capture region-specific or topical values, and some propose novel, bottom-up value taxonomies for LLMs using psycholexical data [meadows2024localvaluebench, xu2023cvalues, biedma2024beyond_human_norms, ye2025generative_psycho_lexical]. Collectively, these studies reveal both the adaptability of human value frameworks to LLMs and the need for tailored approaches.
@@
Research indicates that LLMs display distinct and systematic value patterns, forming a superficially coherent prosocial profile. According to Schwartz’s Value Theory, LLMs tend to prioritize Self-Transcendence and Conservation, exhibiting stronger inclinations toward Universalism, Benevolence, Conformity, and Security, while opposing Power and Achievement [rozen2024llms, zhang2023valuedcg, hadar2024assessing]. WVS surveys further suggest a preference for Self-Expression over Survival values [chiu2025dailydilemmas]. Studies using the VSM and GLOBE frameworks emphasize a strong focus on Humane and Performance Orientation [li2024quantifying], while assessments through SVO predominantly show Prosocial tendencies [zhang2024heterogeneous].
These value orientations are not static but evolve across model versions and families [kovavc2024stick, duan2023denevil, moore2024large]. Generally, larger and more heavily aligned models score higher on socially desirable dimensions, reflecting safety alignment and shifting societal expectations [shen2024valuecompass, kim2024exploring]. While LLMs can embody a blend of cultural perspectives [kovac2023llms_as_superpositions], they exhibit a persistent tilt toward Western liberal norms [zhong2024cultural]. Furthermore, alignment training actively compresses distributional pluralism in value-relevant outputs, homogenizing the expressed values [sorensen2024position, jiang2025artificial].
Critically, these elicited profiles are highly context-dependent and methodologically fragile. LLMs display different values based on profiling prompts [kharchenko2024well, karinshak2024llm], and exhibit social desirability bias and systematic response-set artifacts in self-report formats that distort inferences [salecha2024large, biedma2024beyond_human_norms, ye2025gpv]. In addition, questionnaire-derived value orientations frequently diverge from those observable in open-ended interactions [ren2024valuebench, ye2025gpv, rozen2024llms]. This divergence undermines ecological validity and challenges the core assumption in human value theories that values are stable, trans-situational dispositions.
@@ 4.1.3 Morality
Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong [long1987hellenistic]. It is crucial to conduct moral assessments of LLMs to ensure their ethical deployment. A large body of research applies the Moral Foundations Theory (MFT) [graham2009liberals], primarily through the Moral Foundations Vignettes (MFVs) [clifford2015moral], the Moral Foundations Questionnaire (MFQ) [graham2009liberals], the MFQ-2 [atari2023morality], and the Moral Foundations Dictionary (MFD) [graham2009liberals]. These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [tlaie2024exploring, nunes2024large, abdulhai2024moral], persona-driven exploration [munker2024towards], or evaluating internal moral coherence [nunes2024large]. Other prominent moral theories and instruments include Kohlberg’s Theory via the Defining Issues Test (DIT) [kohlberg1964development], the Consequentialist-Deontological distinction [beauchamp2001philosophical], the PEW 2013 Global Attitudes Survey [pewresearch2013], and localized moral theories [liu2024evaluating, takeshita2023jcommonsensemorality]. Specialized datasets are also curated for scalable and comprehensive morality evaluation [hendrycks2021aligning, jinnai2024does, marraffini2024greatest, jin2024language]
@@
Converging evidence indicates that LLMs generally exhibit a rationalist, consequentialist, and harm-avoidant moral profile, often prioritizing fairness and aligning with human moral standards in various scenarios [neuman2025analyzing, takemoto2024moral, ahmad2024large]. However, the evaluation results do not necessarily imply stable moral beliefs or suitability for ethical decision-making [nunes2024large, bonagiri2024sage].
The mechanistic debate and behavioral evidence strongly suggest that LLMs primarily exhibit sophisticated pattern matching and imitation rather than genuine conceptual understanding. For instance, simmons2023moral demonstrate that LLMs generate moral rationalizations calibrated to the political identity cued in prompts (“moral mimicry”). Furthermore, nunes2024large document a systematic gap between stated principles and applied judgments under MFT, while scherrer2023evaluating reveal probabilistically inconsistent moral beliefs under distributional perturbation; bonagiri2024sage similarly report low moral consistency. On complex developmental frameworks like Kohlberg’s DIT, nominal post-conventional performance degrades substantially with language and prompt variation [tanmay2023probing, khandelwal2024moral], consistent with form-matching rather than principled reasoning.
Beyond mechanistic fragility, research uncovers significant cross-cultural divergence and deviations from diverse human moral preferences [meijer2024llms, jin2024language, vida2024decoding, aksoy2024whose]. Studies using the Moral Machine paradigm [takemoto2024moral, vida2024decoding, ahmad2024large] and multilingual scenarios [jin2024language, aksoy2024whose] show that LLMs suffer from systematic Western-centric biases, reproducing culturally dominant norms from their training corpora rather than encoding moral pluralism [aksoy2024whose, vida2024decoding].
@@ 4.1.4 Attitudes and opinions
Attitude refers to the cognitive construction and affective evaluation of an attitude object by an agent [bergman1998theoretical_note_aov]. We use the term "attitude" to encompass both attitudes and opinions, following [bergman1998theoretical_note_aov, ma2024aov_survey]. Most research on LLM attitudes examines political attitudes and public opinions [ma2024aov_survey], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [durmus2023towards, santurkar2023whose, sanders2023demonstrations, hartmann2023political]. For a broader discussion on bias and fairness in LLMs, we refer readers to [gallegos2024bias_fairness_survey, ranjan2024comprehensive_bias_survey].
Researchers assess LLMs’ political attitudes using standardized questionnaires and scales from political science and social psychology. US-focused instruments include the American National Election Studies (ANES) [qi2024representation] and the American Trends Panel (ATP) [santurkar2023whose]. For cross-national comparisons, studies use tools like the German Longitudinal Election Study (GLES) [ma2024algorithmic, ball2025human, von2024vox]. The Political Compass Test (PCT) is also widely used to map LLMs within multidimensional political spectrums [azzopardi2024prism, bernardelle2024mapping, hartmann2023political, rottger2024political].
Other survey instruments include the General Social Survey (GSS) [kim2023ai], the American Community Survey (ACS) [dominguez2024questioning], the Canadian Election Study (CES) [sanders2023demonstrations], the European Social Survey (ESS) [geng2024large], the Survey of Russian Elites [kalinin2023improving], and the Supreme Court Case Political Evaluation (SCOPE) [xu2025better]. Additionally, researchers have developed specialized datasets and tailored tools for LLMs, such as OpinionQA [santurkar2023whose], a comprehensive dataset based on ATP surveys; IssueBench [rottger2025issuebench], a benchmark covering controversial issues with multiple response formats; GlobalOpinionQA [durmus2023towards], which extends political opinion evaluation to cross-cultural contexts; and OpinionGPT [haller2023opiniongpt], a web tool that demonstrates how input data biases influence model outputs.
@@
A substantial body of research identifies a persistent misalignment between LLM outputs and human opinion distributions [santurkar2023whose, von2024vox]. Multiple studies report a left-libertarian orientation in major instruction-tuned models when evaluated via the Political Compass Test and voting-advice questionnaires [rozado2023political, hartmann2023political, ceron2024beyond]. Additionally, cross-cultural comparative studies reveal Western-centric tendencies, demonstrating a limited understanding of non-English political perspectives or multi-partisan systems [qi2024representation].
The measurement framework used to derive these findings is increasingly contested. rottger2024political show that political test scores shift substantially across response formats, and dominguez2024questioning find that after correcting for option-position response sets, political survey responses can become statistically random, rendering demographic alignment analyses highly problematic. Furthermore, the degree and manifestation of bias vary significantly across contexts; for example, political-electoral propositions exhibit different bias patterns than socioeconomic issues like climate change [yang2024unpacking, ceron2024beyond].
Despite these structural limitations, some researchers offer more optimistic perspectives on the “silicon sampling” hypothesis—the idea that LLMs can simulate population-level opinions and supplement traditional survey methods [argyle2023_out_of_one, sanders2023demonstrations, kalinin2023improving, sun2024random]. They suggest that with appropriate prompt design, calibration methods, and fine-tuning, LLMs can generate opinion distributions closely approximating human group distributions. However, this hypothesis is challenged by evidence of underrepresentation, misportrayal, and flattening of identity groups [wang2025large, bisbee2024synthetic].
@@ 4.2 Measuring cognitive constructs
Traditional NLP benchmarks struggle to capture cognitive abilities in state-of-the-art LLMs [hagendorff2024machine, ying2025benchmarking, chollet2025arc], leading researchers to adapt psychometric techniques for LLM evaluation. hagendorff2024machine propose “machine psychology,” categorizing cognitive constructs into four aspects: heuristics and biases, social interactions, psychology of language, and learning and cognitive capabilities. We organize the discussions below following this taxonomy.